Nevey:

A sequence has its first term equal to 4, and each term of the sequence is obtained by adding 2 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 2 and f(n) = f(n − 1) + 4; n > 1 f(1) = 4 and f(n) = f(n − 1) + 2n; n > 1 f(1) = 2 and f(n) = f(n − 1) + 4n; n > 1 f(1) = 4 and f(n) = f(n − 1) + 2; n > 1

Nevey:

@Kamauri

Kamauri:

We can describe the sequences two ways: One: T(1) = 4 and T(n+1) = T(n) + 2 Two: T(n) = 2n + 2 You excluded the following so I did not know which is the correct version in this case (both convey the same point and mean the same thing)

Nevey:

that explains nothing to me .-.

Kamauri:

i know im sorry im stuck on this one

Nevey:

its alright.

Nevey:

@BenLindquist @Razor

dude:

Arithmetic sequences are written as \(a_n=a_1+(n-1)d\) \(a_1\) is the first term d is the "rate"